How many ways can you draw a nine or a club from a deck of cards? And why do you subtract it by 1??
Answer by richard1234(7193)

# of 9's: 4
# of clubs: 13

# of 9's + # of clubs = 4+13 = 17

But you overcounted the 9 of clubs (we counted it with 9's and with clubs), so we have to subtract 1 (since we counted it twice but only want it once), so the answer is 4+13-1 = 16.